ananazzaa:

Solve the following system of equations and show all work. y = 2x^2 y = −3x − 1

Axelotle:

ok so you need to get the variable alone, so for the first problem what is the opposite of the ^ and for the second problem what is the opposite of a subtraction sign??

ananazzaa:

opposite of subtraction is addition

ananazzaa:

anddd the opposite of ^ isnt that division

Axelotle:

bingo sofor the second problem add 1 to 1 and cross it out bc that is practically a negative and a positive, and then for the 1st problem also yes so divide

ananazzaa:

so for the 2nd y+1=-3x

Axelotle:

yes there u go, now try and do the other one! :}

ananazzaa:

wait cant i just do 2x^2+3x+1=0 then just factor or use the quadratic formula

Axelotle:

yep there u go!! :]

ananazzaa:

so i used the quadratic formula

ananazzaa:

i got that x = -1/2 and -1

Axelotle:

yep there u go :D

ananazzaa:

yay